OpenStudy (anonymous):

if 51% = $25,000, what percent = $20,000?

OpenStudy (anonymous):

51% of a value, X, = $25,000

OpenStudy (anonymous):

\[.51 * X = 25,000\]

OpenStudy (anonymous):

solve for X \[X = 25,000/(.51)\]

OpenStudy (anonymous):

then you can solve for "what percent of X equals $20,000" ?

OpenStudy (anonymous):

which would be \[(some percentage) * X = 20,000\]

OpenStudy (anonymous):

\[ X = 25,000/(.51)\] so \[(percentage) * {25,000/(.51)} = 20,000 \] so \[(percentage) = 20,000 / [25,000/(.51)]\]

OpenStudy (anonymous):

hope this helps

OpenStudy (anonymous):

So, what % = 20,000?

OpenStudy (anonymous):

you can easily solve that following the steps I laid out ;)

OpenStudy (anonymous):

you could also set up a ratio and cross multiply:\[\left( .51\div25k \right)=\left( x \div20k \right)\] solve for x.